Bulla Recording, Inc., wishes to maintain a growth rate of 12 percent per year and a debt–equity ratio of .40. Profit margin is 5.3 percent, and the ratio of total equity to sales is constant at 53.59 percent. Is this growth rate possible? To answer, determine what the dividend payout ratio must be. How do you interpret the result?
Answer #1

Given data,

Growth rate (g) = 12%

debt-equity ratio(D/E) = 0.40

Profit margin = 5.3%

Total equity to sales = 53.59%

As we know, Growth rate (g) = Plowback ratio * Return on equity

And Plowback ratio = 1-dividend payout ratio

ROE = Net income / Equity

So, g = (1-d)*(NI/E)

or g = (1-d)*(NI/sales)*(sales/E) multiplying and dividing by sales

Putting NI/sales and E/sales in the above equation we get

g = (1-d)*(0.053/0.5359) => g=(1-d)*0.099

as we know, normally dividend payout ratio can be between 0% to 100%,

So, when d=0, g = 9.9%

and when d=1, g= 0%

so given the scenario, g is always less than 9.9%.

So, Bulla Recording, Inc. cannot maintain an growth rate of 12% with above given data.